OpenStudy (anonymous):

Factor please: 8x^2-6x-9=0

OpenStudy (anonymous):

need to know how to do it and why please

OpenStudy (misssunshinexxoxo):

8x^2 - 6x - 9 = 0 ac = 8(-9) = - 72 Factors of - 72 that add to - 6 are +6 & - 12 8x^2 + 6x - 12x - 9 = 0 2x(4x + 3) - 3(4x + 3) = 0 (2x - 3)(4x + 3) = 0 x = 3/2, - 3/4

OpenStudy (anonymous):

hang on lemme process this

OpenStudy (anonymous):

Why is it (2x-3) (4x+3)=0 did u skip a step?

OpenStudy (misssunshinexxoxo):

No that is the process after factoring

OpenStudy (misssunshinexxoxo):

Plug in the numbers, factor, and you receive the solution